# MultiScale Contextual Bandits for Long Term Objectives

### v1

@@ Abstract
The feedback that AI systems (e.g., recommender systems, chatbots) collect from user interactions is a crucial source of training data. While short-term feedback (e.g., clicks, engagement) is widely used for training, there is ample evidence that optimizing short-term feedback does not necessarily achieve the desired long-term objectives. Unfortunately, directly optimizing for long-term objectives is challenging, and we identify the disconnect in the timescales of short-term interventions (e.g., rankings) and the long-term feedback (e.g., user retention) as one of the key obstacles. To overcome this disconnect, we introduce the framework of MultiScale Policy Learning to contextually reconcile that AI systems need to act and optimize feedback at multiple interdependent timescales. For any two levels, our formulation selects the shorter-term objective at the next lower scale to optimize the longer-term objective at the next higher scale. As a result, the policies at all levels effectively optimize for the long-term. We instantiate the framework with MultiScale Off-Policy Bandit Learning (MSBL) and demonstrate its effectiveness on three tasks relating to recommender systems and text generation.
Machine Learning, ICML
[…]
[ backgroundcolor=orange!10, linecolor=orange!50, roundcorner=5pt, linewidth=1pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=10pt, innerbottommargin=10pt ]custombox
@@ 1 Introduction
Many interactive AI systems (e.g., recommender systems, conversational systems) use abundantly collected short-term feedback for learning. However, it is well known that over-optimization of short-term feedback can adversely affect the long-term goals (Mansoury et al., 2020; Milli et al., 2021). Similarly, it can lead to issues like reward hacking (Skalse et al., 2022; Pan et al., 2022, 2024) and user manipulation (Kasirzadeh & Evans, 2023; Everitt et al., 2021). For example, optimizing for engagement on social media platforms can lead to clickbait-y or toxic feeds. This neither reflects the user preferences nor the platform’s goals of retaining users, and it raises the question of what objectives to align to in the first place (Stray et al., 2021). We wish to design systems that optimize for long-term objectives, where these objectives are beneficial outcomes for various stakeholders in the system (Abdollahpouri et al., 2019).
A key problem in achieving this goal is that the long-term feedback (e.g., user retention) is at a different timescale than the short-term interventions (e.g., rankings of recommended products) that are used to optimize them.
@@
For instance, optimizing rankings for clicks in a recommender system is relatively straightforward since clicks can be attributed to an individual ranking. Indeed, learning methods such as bandits are widely used to learn ranking policies (Li et al., 2011). Now, consider optimizing rankings for the feedback on users’ subscription renewal, which is observed monthly. We note the disconnect between the timescale of rankings presented and the long-term objective of subscription renewal, making this a much harder problem than the previous one. While a flat MDP approach in principle is able to model sequential dependencies to achieve long-term goals, in practice the large resulting state spaces, the associated credit-assignment problems, and the sparsity of long-term feedback prohibit straightforward applications of reinforcement learning (Maystre et al., 2023).
To overcome this problem, our approach is to contextually reconcile the disconnect between short-term and long-term objectives by learning interventions and policies at multiple timescales. To this end, we propose a new framework – MultiScale Policy Learning – that involves multi-level formulation for optimizing long-term objectives. For every macro (upper level operating at a slower timescale) and micro (lower level operating at a faster timescale) level, we specify an interface such that the shorter-term objective is selected to optimize the longer-term objective. We demonstrate the generality of this framework with different interface designs. To instantiate policy learning at multiple levels, we introduce MultiScale Off-Policy Bandit Learning (MSBL) as a simple and practical recursive Algorithm. Our method learns a sequence of nested policies that enable the use of both abundant short-term and sparse long-term data for optimizing long-term outcomes. We demonstrate the effectiveness of our approach empirically on three tasks ranging from recommender to conversational systems.
2 Motivation
@@
To motivate our approach, consider a recommender system for a video streaming platform as depicted in Figure  1. At the lowest level, the system acts by providing rankings of recommended products. These rankings (short-term interventions) can be optimized for clicks (short-term reward). While clicks are an important signal for learning that is abundantly available at the lowest level of Figure  1, an unmitigated maximization of clicks is not necessarily aligned with higher-level goals. In particular, the platform may sacrifice some clicks, if that leads to a higher weekly return rate. This metric is a more reliable indicator of user satisfaction, but it is available at a slower timescale. Finally, at the highest level, the platform ultimately cares about user retention and subscription renewal, since this is their economic basis. This feedback lives at an even slower timescale, and it is thus even more scarce but even more valuable.
As a second example, consider a GenAI tutoring system. This system will want to use abundant chat-engagement feedback at the lowest level, but it is eventually judged by how well it achieves the learning outcomes of a course at the highest level – with potentially multiple levels of intermediate metrics in between. Similar multi-scale levels of feedback metrics exist also for other AI systems, and additional settings are discussed in Appendix A.
We make the following observations. First, the feedback observed at multiple levels exists as concepts at that level and can not be observed at lower levels. For example, weekly return rate does not exist at the level of individual rankings, as it is affected by potentially many rankings. Second, we argue that longer-term feedback is best evaluated in the context of longer-range interventions, and that each feedback level should come with its own interventions. Examples of longer-term interventions can be diversity boosts, aggressiveness of spam filtering, the decoding strategy of an LLM policy, etc. This raises the question of how to leverage the interventions and feedback at different timescales for using the short-term and long-term data to effectively optimize for long term outcomes.
@@ 3 Related Works
Hierarchical Reinforcement Learning. Hierarchical RL (HRL) approaches, such as the options framework (Sutton et al., 1999; Bacon et al., 2017) and feudal learning (Dayan & Hinton, 1992), learn to operate on different levels of temporal abstraction. HRL aims to accomplish complex tasks by dividing them into sub-goals with the higher level assigning sub-tasks to the next lower level, such as in robotic applications (Pateria et al., 2021). While our framework shares the central idea of elevating different timescale feedback to separate levels, there are several differences. We study the tradeoff in longer-term/shorter-term feedback in contrast to the clear subgoal decomposition of HRL, in which the discovery of the subgoals and their execution is critical and guides the micro policy. Our approach, on the other hand, leverages the reward structure with shorter-term dependence on longer term feedback signals. Similarly, the macro actions (options) in HRL are abstractions over micro actions and the goal is to combine these options for arbitrary tasks. In contrast, the macro interventions (e.g., decoding strategy, preference weights etc.) we consider interface with the micro policy to steer towards improving long-term outcomes, even at the expense of shorter-term rewards. Our approach is motivated by reconciling the disconnect in actions and feedback by having a one-to-one correspondence between them at each timescale.
Long Term Optimization. There has been a growing interest in the study of recommender systems that go beyond optimizing engagement and clicks (McDonald et al., 2023; Brantley et al., 2024; Maystre et al., 2023; Milli et al., 2021; Cai et al., 2023; Agarwal et al., 2024). (Maystre et al., 2023) provide an RL perspective to long-term learning in recommender systems and discuss challenges with credit attribution. (Brantley et al., 2024) formulate the problem with long-term metric as a constraint satisfaction with minimum impact on short-term engagement metric. These works operate on a single timescale and assume that the macro intervention is given and fixed. Our contribution is to learn the macro interventions themselves from long-term feedback and introduce learning policies and interventions for different timescales. In fact, these methods can be used for a single level-specific learning within our MultiScale framework, making them a special case of our approach.
Multi-Objective Optimization. Recent work has highlighted the importance of selecting weights for linear scalarization of multi-objective learning in recommender systems (Milli et al., 2024; Jeunen et al., 2024; Wilm et al., 2024) and in text generation tasks (Yang et al., 2024). Our framework achieves multi-objective learning by learning a family of policies at the micro level and a separate policy at the macro level. As a result, unlike prior works, we propose deploying a sequence of policies at inference. Separately, while the idea of family of policies has been explored previously in (Dosovitskiy & Djolonga, 2020; Wilm et al., 2024; Yang et al., 2024), these works do not learn the preference weights, operate on a single level and are not related to learning long-term outcomes. We learn for the first time a sequence of nested contextual bandit policies for long term optimization. Additional related works are deferred to the Appendix B.
@@ 4 Multi-Scale Policy Framework
Our key insight is to reconcile the tradeoff between longer-term and shorter-term feedback by leveraging interventions and feedback in the system at multiple levels. Extending the conventional contextual-bandit approach, we first define how multiple scales of contexts, actions, and rewards can be tied together through a nested set of policies. This results in a new modeling framework where data-rich lower levels can handle large action spaces, while upper levels that receive increasingly sparse feedback can be conservative in the size of their action spaces. We then show how to learn the policies at each level to eventually optimize the most long-term reward.
@@
We begin by defining how contexts, actions, and rewards at multiple levels interact through a nested set of policies. For simplicity, we restrict our notation to two levels – the macro level operating at the slower timescale $t{2}$ and the micro level operating at the faster timescale $t1$. As we will see later, the framework naturally extends to multiple scales of policy learning.
Figure 2: (a) At inference, a macro action indexes to select the particular micro policy from a family of micro policies. The macro action space $\mathcal{A}^{L2}$ is isomorphic to the family of policies $\Pi^{L1}$ (b) Learning micro policies: Abundant micro-level data is used to learn a family of promising policies $\Pi^{L1}$ using policy or feedback modification (c) Macro-level data is used to learn a macro policy. For more than two levels, (b) and (c) are recursively called narrowing down micro policy space/ macro action space.
@@
Multi-Scale Contexts. At the upper level, contexts $x^{L2}\sim p(x^{L2})$ arrive at timescale $t{2}$. An example of an upper-level context could be a user as described by demographic features or some long-term profile. For each such upper-level context, a sequence of lower-level contexts $x^{L1}\sim p(x^{L1}|x^{L2})$ is drawn conditionally on $x^{L2}$. Such lower-level contexts could be search queries, chat requests, or opportunities for push notifications.
Multi-Scale Policies. We consider nested policies across both levels, and each policy can be stochastic. For each upper-level context $x^{L2}$, the upper-level policy $\pi^{L2}$ selects an action $a^{L2}$ from action space $\mathcal{A}^{L2}$.
$a^{L2}\sim\pi^{L2}(a^{L2}|x^{L2})$
@@
Importantly, as shown in Figure 2 (a), the action $a^{L2}$ corresponds to a lower-level policy $\pi^{L1}_{a^{L2}}$, which means that the action space $\mathcal{A}^{L2}$ at the upper level is isomorphic to a family of policies $\Pi^{L1}=\{\pi^{L1}_{a^{L2}}:a^{L2}\in\mathcal{A}^{L2}\}$ at the lower level.
$\mathcal{A}^{L2}\cong\Pi^{L1}$
@@
Multi-Scale Rewards. Similar to contexts and actions, we also observe rewards at different timescales. At the lower level, we observe a reward $r^{L1}\sim p(r^{L1}|x^{L1},a^{L1})$ for each lower-level action $a^{L1}$ and context $x^{L1}$. This corresponds to the short-term engagement (e.g., clicks) optimized in conventional approaches. A key difference is that we also record rewards $r^{L2}\sim p(r^{L2}|x^{L2},\pi^{L1}_{a^{L2}})$ for each upper-level action $a^{L2}$ and context $x^{L2}$, which corresponds to a long-term metric like weekly returns or subscription renewals.

### v2

@@ Abstract
The feedback that AI systems (e.g., recommender systems, chatbots) collect from user interactions is a crucial source of training data. While short-term feedback (e.g., clicks, engagement) is widely used for training, there is ample evidence that optimizing short-term feedback does not necessarily achieve the desired long-term objectives. Unfortunately, directly optimizing for long-term objectives is challenging, and we identify the disconnect in the timescales of short-term interventions (e.g., rankings) and the long-term feedback (e.g., user retention) as one of the key obstacles. To overcome this disconnect, we introduce the framework of MultiScale Policy Learning to contextually reconcile that AI systems need to act and optimize feedback at multiple interdependent timescales. Following a PAC-Bayes motivation, we show how the lower timescales with more plentiful data can provide a data-dependent hierarchical prior for faster learning at higher scales, where data is more scarce. As a result, the policies at all levels effectively optimize for the long-term. We instantiate the framework with MultiScale Off-Policy Bandit Learning (MSBL) and demonstrate its effectiveness on three tasks relating to recommender and conversational systems.
@@ 1 Introduction
Many interactive AI systems (e.g., recommender systems, conversational systems) use abundantly collected short-term feedback for learning. However, it is well known that over-optimization of short-term feedback can adversely affect the long-term goals [18, 23]. Similarly, it can lead to issues like reward hacking [30, 25, 26] and user manipulation [15, 10]. For example, optimizing for engagement on social media platforms can lead to clickbait-y or toxic feeds. This neither reflects the user preferences nor the platform’s goals of retaining users. We wish to design systems that optimize for long-term objectives, that are beneficial for various stakeholders in the system [1].
A key problem in achieving this goal is that the long-term feedback (e.g., user retention) is at a different timescale than the short-term interventions (e.g., rankings of recommended products) that are used to optimize them. For instance, optimizing rankings for clicks in a recommender system is relatively straightforward since clicks can be attributed to an individual ranking. Indeed, learning methods such as bandits that optimize for users’ immediate response to recommendations such as clicks, likes, views, etc are widely used to learn ranking policies [16, 39]. Now, consider optimizing rankings for the feedback on users’ subscription renewal, which is observed monthly. We note the disconnect between the timescale of rankings presented and the long-term objective of subscription renewal, making this a much harder problem than the previous one. While a Markov Decision Process can in principle model sequential dependencies to achieve long-term goals, in practice the large resulting state spaces, credit-assignment problems, and the sparsity of long-term feedback prohibit straightforward applications of reinforcement learning [20].
To overcome this problem, our approach is to contextually reconcile the disconnect between short-term and long-term objectives by learning interventions and policies at multiple timescales.
@@
Consider a recommender system for a video streaming platform as depicted in Figure  1. At the lowest level, the system acts by providing rankings of recommended products. These rankings (short-term interventions) can be optimized for clicks (short-term reward). While clicks are an important signal for learning that is abundantly available at the lowest level of Figure  1, an unmitigated maximization of clicks is not necessarily aligned with higher-level goals. In particular, the platform may sacrifice some clicks if that leads to a higher weekly return rate. This metric is a more reliable indicator of user satisfaction, but it is available at a slower timescale. Finally, at the highest level, the platform ultimately cares about user retention and subscription renewal. This feedback lives at an even slower timescale, and it is thus more scarce but even more valuable. Similar multi-scale levels of feedback metrics also exist for other AI systems (e.g., tutoring chatbots that aim to achieve long-term learning outcomes), and additional settings are discussed in Appendix B.
The main contributions of this work are as follows:
  1. 1.
    We introduce a new framework – MultiScale Policy Learning – that formalizes a multi-level approach for optimizing long-term objectives. The framework introduces a recursive construction of priors that are informed by data at the lower levels to speed up learning at the higher levels.
  2. 2.
    We demonstrate the practicality and generality of this framework by developing MultiScale Off-Policy Bandit Learning (MSBL) as a simple recursive algorithm for training multi-scale contextual bandit policies. It includes two widely applicable constructions for nested training of policies that enable the use of both abundant short-term and sparse long-term data for optimizing long-term outcomes.
  3. 3.
    We demonstrate the effectiveness of our approach empirically on three tasks ranging from recommender to conversational systems. Ablations show robustness of our method for optimizing the long term reward.
@@ 2 Related Works
Hierarchical Reinforcement Learning. Hierarchical RL (HRL) approaches, such as the options framework [32, 3] and feudal learning [7], learn to operate on different levels of temporal abstraction. HRL aims to accomplish complex tasks by dividing them into sub-goals with the higher level assigning sub-tasks to the next lower level, such as in robotic applications [27]. While our framework shares the idea of creating a hierarchy, the type of hierarchy is fundamentally different. In particular, our tasks do not have a subgoal decomposition as in HRL, where the discovery of the subgoals and their execution is critical and guides the micro policy. Instead, we exploit the hierarchy of feedback timescales to construct a hierarchical prior over the policy space to speed up learning for sparse long-term feedback. Our goal is to steer towards improving long-term outcomes, even at the expense of shorter-term rewards. This is fundamentally different from the macro actions (options) in HRL, which are abstractions over micro actions and the goal is to combine these options as subroutines.
Long Term Optimization. There has been a growing interest in the study of recommender systems that go beyond optimizing engagement and clicks [22, 5, 20, 23, 6, 2]. Maystre et al. [20] provide an RL perspective to long-term learning in recommender systems and discuss challenges with credit attribution. Other approaches, such as [5] do not learn at the macro level and instead formulate long-term macro interventions that they aim to fulfill with minimum impact on short-term engagement metrics. These works operate on a single timescale and assume that the macro intervention is given and fixed. Our work can be viewed as learning the macro interventions themselves from long-term feedback. In fact, these methods can be used for a single level-specific learning within our MultiScale framework, making them a special case of our approach.
Multi-Objective Optimization. Recent work has highlighted the importance of selecting weights for linear scalarization for multi-objective learning in recommender systems [24, 13, 35] and in text generation tasks [36]. A different line of work explores learning conditional policies as a single family of policies [8, 35, 36]. Our work leverages the idea of conditional policies and multi-objective learning from the perspective of optimizing for long-term outcomes. In doing so, we elevate single-stage learning to multiple levels with the macro level contextually selecting the objective to optimize at the micro level.
Additional related works are deferred to Appendix C.
@@ 3 Multi-Scale Policy Framework
We begin by providing a PAC-Bayesian motivation for why a hierarchical approach that exploits feedback across multiple scales can be substantially more data efficient. For simplicity of notation, we restrict to two levels – the micro level operating at the faster timescale $t1$ and the macro level operating at the slower timescale $t{2}$. As we will see later, the framework naturally extends to multiple scales of policy learning.
Our goal is to learn a policy $\pi(a|x)$ that selects an action $a$ for a given context $x$. We assume that contexts are drawn i.i.d. from an unknown distribution $x\sim p(x)$, but we conjecture that our approach can be extended to stateful models as well. Our approach rests on the realization that we observe rewards at different timescales in many AI systems. For two levels, we observe a reward
$r^{L1}\sim p(r^{L1}|x,a)$
at the micro level for each action $a$ and context $x$, corresponding to the short-term engagement (e.g., clicks). We also record rewards
$r^{L2}\sim p(r^{L2}|(x_{i},a_{i}),\ldots(x_{i+T},a_{i+T}))$
after we have taken a sequence of actions. This corresponds to the long-term feedback (e.g., weekly returns, subscription renewal) which we would like to optimize, since it typically better reflects stakeholder objectives. This leads us to the following policy learning objective,
$\pi^{L2*}\leftarrow\operatorname*{\arg\max}_{\pi\in\Pi}V^{L2}(\pi),~{}~{}\mbox{ where }~{}V^{L2}(\pi)=\mathbb{E}_{x,a,r^{L2}}[r^{L2}].$ (1)
Unfortunately, for a large and complex policy space $\Pi$, finding the optimal policy $\pi^{*}$ by simply replacing the expected reward $V^{L2}(\pi)$ with its empirical estimate $\hat{V}^{L2}(\pi)$ on some training data $D^{L2}$ is typically intractable. The long-term reward is too infrequent for learning policies from scratch. As a result, existing approaches predominantly [37] optimize the more frequent reward signal $r^{L1}$ at the micro level.
$\pi^{L1*}\leftarrow\operatorname*{\arg\max}_{\pi\in\Pi}V^{L1}(\pi),~{}~{}\mbox{ where }~{}V^{L1}(\pi)=\mathbb{E}_{x,a,r^{L1}}[r^{L1}].$ (2)
Note that even the policy $\pi^{L1*}$ that perfectly optimizes the micro level reward can have substantially worse reward at the macro level, $V^{L2}(\pi^{L1*})<V^{L2}(\pi^{L2*})$. However, $\pi^{L1*}$ is typically much better than a random policy from $\Pi$. This raises the following question.
How do we exploit feedback at the micro level to learn faster at the macro level?
To provide a theoretical motivation, we make the following PAC-Bayesian argument. PAC Bayes generalization bounds [17, 21] provide uniform convergence over all posterior distributions $Q(\pi)$ for any given prior distribution $P(\pi)$. Specifically, with probability $1-\delta$,
$\big{|}\mathbb{E}_{\pi\sim Q}[V^{L2}(\pi)]-\mathbb{E}_{\pi\sim Q}[\hat{V}^{L2}(\pi;D)]\big{|}\leq O\bigg{(}\sqrt{\frac{KL(Q||P)+ln(1/\delta)}{n}}\bigg{)}.$ (3)
Since this bound holds for all $Q$, it also holds for any (approximately) optimal posterior $Q^{L2*}$ that maximizes the macro level reward. For a discrete policy space, $Q^{L2*}$ is the Dirac delta distribution centered on $\pi^{L2*}$. The bound states that this learning problem is ‘easy’ (i.e., requires a small number $n$ of training examples) if the KL-divergence between $Q^{L2*}$ and the prior $P$ is small.
Can we improve the prior for the macro level with data from the micro level?
While the optimal policy $\pi^{L1*}$ at the micro level may be suboptimal at the macro level, a policy that is learned based on finite data at the micro level via $\hat{\pi}^{L1}\leftarrow\operatorname*{\arg\max}_{\pi\in\Pi}\hat{V}^{L1}(\pi)$ can provide useful prior information for learning at the macro level.
Consider the following illustrative example, where each policy $\pi(.|.,\theta)\in\Pi$ is defined via a parameter vector $\theta$. We denote $\theta^{L1}$ as the parameters of policy $\hat{\pi}^{L1}$, and $\theta^{L2*}$ as those of the optimal macro policy $\pi^{L2*}$. For simplicity of demonstration, we define the target policy distribution as $Q^{L2*}=N(\theta^{L2*},\Sigma^{L2})$, an uninformed prior distribution $P_{0}=N(\theta_{0},\Sigma_{0})$ for some arbitrary $\theta_{0}$, and an informed prior $P^{L1}=N(\theta^{L1},\Sigma^{L1})$ centered at the learned micro policy $\hat{\pi}^{L1}$. The difference in training samples $n_{0}-n^{L2}$ to get the same confidence interval for $Q^{L2*}$ in Equation (3) is proportional to $KL(Q^{L2*}||P_{0})-KL(Q^{L2*}||P^{L1})$. We show in Section E.1 that for an appropriately chosen $\Sigma^{L1}$, the improvement in required training samples by moving to the informed prior $P^{L1}$ is at least
$n_{0}-n^{L2}\propto KL(Q^{L2*}||P_{0})-KL(Q^{L2*}||P^{L1})\in O(|\theta^{L2*}-\theta_{0}|_{M}-|\theta^{L2*}-\theta^{L1}|_{M})$ (4)
where $|\theta^{L2*}-\theta_{0}|_{M}$ is the squared Mahalanobis distance in the parameter space $\theta$.
This shows that the policy $\hat{\pi}^{L1}\sim P^{L1}$ learned at the micro level can provide training sample savings at the macro level, if $\hat{\pi}^{L1}$ and $\pi^{L2*}$ are close compared to the uninformed prior $P_{0}$. In particular, if we can learn almost all parameters of $\pi^{L2*}$ at the micro level, the distance $|\theta^{L2*}-\theta^{L1}|_{M}$ will be small, resulting in higher macro level sample savings. The figure on the right illustrates how an informed prior $P^{L1}$ pulls the center of the prior $P_{0}$ closer to the parameters of $\pi^{L2*}\sim Q^{L2*}$. To quantify the reduction in macro level training data, we construct a numerical example (detailed in Section E.1) with $\theta\in\operatorname{\mathbb{R}}^{50}$, such that 49 parameters are learned well at the micro level and only 1 parameter needs to be adjusted at the macro level. This results in saving $\approx 98\%$ training samples [^1] This calculation does not consider the investment of training samples $n^{L1}$ for constructing $P^{L1}$. Generally these $n^{L1}$ samples are significantly cheaper than the macro samples $n^{L2}$, since they are $T$ times more frequent., which means gathering enough macro level data may only take weeks instead of years.
4 Multi-Scale Policy Learning
In order to put the theoretical insights from the previous section into a practical policy learning algorithm, we introduce a factorization of contexts and policies at each level. We propose the policy factorization in a way that the micro level learns a large part of the parameter space, even if it is not aligned with the long term expected reward. This simplifies learning at macro level compared to learning from scratch.
@@ 4 Multi-Scale Policy Learning
Multi-Scale Contexts. The context $x\sim p(x)$ can be factorized as $p(x)\triangleq p(x^{L2})\cdot p(x^{L1}|x^{L2})$. At the upper level, contexts $x^{L2}\sim p(x^{L2})$ arrive at timescale $t{2}$. An example of an upper-level context could be a user as described by demographic features or some long-term profile. For each such upper-level context, a sequence of lower-level contexts $x^{L1}\sim p(x^{L1}|x^{L2})$ is drawn conditionally on $x^{L2}$. Such lower-level contexts could be search queries, or chat requests.
Multi-Scale Policies. We consider the following factorization of policy space $\Pi$,
$\Pi\triangleq\Pi^{L1}\cdot\Pi^{L2},$
where $\Pi^{L1}$ consists of micro policies, that as we will see later, can provide a strong inductive bias for the long term optimal policy. In contrast, the macro level is only concerned with learning within the space of reasonable policies obtained after micro learning and has a much smaller policy space $\Pi^{L2}$.
Specifically, for each upper-level context $x^{L2}$, the upper-level policy $\pi^{L2}$ selects an action $a^{L2}$ from action space $\mathcal{A}^{L2}$. Examples of $a^{L2}$ are diversity boosts, aggressiveness of spam filtering, the decoding strategy of an LLM policy, etc.
$a^{L2}\sim\pi^{L2}(a^{L2}|x^{L2})~{}~{}~{};~{}~{}~{}\pi^{L2}\in\Pi^{L2}$
@@
Importantly, as shown in Figure 2 (a), the action $a^{L2}$ corresponds to a lower-level policy $\pi^{L1}_{a^{L2}}$. This means that the action space $\mathcal{A}^{L2}$ at the upper level is isomorphic to a family of policies $\hat{\Pi}^{{L1}}=\{\hat{\pi}^{L1}_{a^{L2}}:a^{L2}\in\mathcal{A}^{L2}\}$, learned empirically at the lower level.
$\mathcal{A}^{L2}\cong\hat{\Pi}^{{L1}}$